OpenStudy (anonymous):

Help Indicate the equation of the line that is the perpendicular bisector of the segment with endpoints (4, 1) and (2, -5).

OpenStudy (anonymous):

Find the gradient between the 2 points: m = y2-y1/x2-x1 = 1 - (-5)/4-2 = 3 The perpendicular gradient will be -1/3 as 3*(-1/3) = -1 Find the midpoint between the 2 points: \[= (\frac{ x2+x1 }{2 } + \frac{ y2+y1 }{ 2 })\] \[= (\frac{ 4+2 }{2 } + \frac{1+(-5) }{ 2 })\] = (3, -2) The equation is given by y-b=m(x-a), substituting in the midpoint y+2 = -1/3(x-3) 3y+6 = -x+3 3y = -x-3

OpenStudy (anonymous):

Thank youu

OpenStudy (anonymous):

No problem :) I believe that its correct but its really late here xD